# Radiopurity screening of materials for rare event searches by neutron activation at the TRIGA reactor of Pavia

### v1

@@ Abstract
In the framework of physics experiments searching for rare events, the selection of extremely radiopure materials is a challenging task, as the signal of interest is often hidden by instrumental background. Neutron activation is a powerful technique to measure trace contaminants with high sensitivity but, to be properly applied, it requires a good characterization of neutron irradiation and $\gamma$-spectroscopy facilities. This paper presents the state-of-the-art workflow adopted by the Radioactivity Laboratory of the University of Milano-Bicocca for radiopurity screening of materials by neutron activation performed at the TRIGA reactor in Pavia. The ultimate sensitivity of the described workflow, in the absence of interfering activation products and without the application of radiochemical and/or active background rejection techniques, is $<10^{-14}$ g/g for 40K, $<10^{-12}$ g/g for 238U and of the order of $10^{-12}$ g/g for 232Th contaminations. Further details are here provided to address systematic uncertainties related to neutron irradiation that may bias results. To this aim, a dedicated neutron activation campaign was performed and the data were analyzed exploiting a Monte Carlo simulation model of the reactor and applying an unfolding technique to obtain a comprehensive characterization of the neutron flux in typical irradiation configurations. The results of this work provide a valuable benchmark for the application of neutron activation in future radiopurity screening campaigns.
pacs:
PACS-keydiscribing text of that key and PACS-keydiscribing text of that key
@@ 1 Introduction
Experiments searching for rare particle physics events demand increasingly higher levels of radiopurity for the materials that make up the detectors heusser:1995 . The ultimate experimental sensitivity critically depends on the level of residual radioactive background in the energy region of interest, which could hamper the detection of the very faint signals being sought. Besides locating these experiments in deep underground laboratories to reduce cosmogenic-induced events Laubenstein:2004 , surrounding the detectors with massive shielding to attenuate environmental background Smith:2008 , and employing particle identification methods to discard spurious energy depositions Budjas:2009zu , it is crucial to construct the detector using materials with the lowest possible concentration of radioactive contaminants. In this respect, the assay techniques also require increasingly higher levels of sensitivity to evaluate the radiopurity of materials.
Typical contaminants include primordial radionuclides such as 40K, 232Th, 235U and 238U, as well as their decay chain daughters (for Th and U isotopes). Depending on the materials, a few cosmogenic and artificial long-lived radionuclides, such as 60Co and 137Cs, may also be of concern. The most widely used screening technique to study bulk contamination of materials is $\gamma$-spectroscopy Laubenstein:2017yjj , which has the significant advantage of being sensitive to most radionuclides in natural decay chains. Alpha spectroscopy is typically used to evaluate the presence of 232Th, 238U, and their $\alpha$-decaying daughters (particularly radon progenies) on material surfaces Bunker:2020sxw . Complementary bulk contamination screening techniques include mass concentration measurement methods like Inductively Coupled Plasma Mass Spectrometry (ICP-MS) and Neutron Activation Analysis (NAA) Arpesella:2001iz ; Nisi:2009 . Although these techniques are limited to detecting only the parent nuclides, such as 232Th and 238U, without the ability to assess the concentrations of their daughters, they can achieve much higher analytical sensitivities compared to $\gamma$-spectroscopy.
@@ 2 Experimental methodology
The main step of NAA is neutron irradiation. The contaminants to be measured are targets of neutron-induced reactions, usually at a research reactor, which lead to the production of radionuclides. Subsequently, $\gamma$-spectroscopy measurements are performed to measure the induced activities and determine the concentrations of contaminants in the sample. High-Purity Germanium (HPGe) detectors are the most widely used $\gamma$-spectrometers, because of their excellent energy resolution.
@@
[…]
@@
[…]
@@
where the superscript $std$ is used to label the parameters of the standard. It is worth noting here that if the standard has the same geometry and materials of the sample, the efficiency ratio $\varepsilon_{\gamma}/\varepsilon_{\gamma}^{std}$ can be assumed equal to 1, otherwise a Monte Carlo simulation is usually needed to determine this corrective factor with good accuracy.
@@
[…]
@@ 2.1 NAA sensitivity
After neutron irradiation, the samples are extracted from the reactor and prepared for $\gamma$-spectroscopy, typically by means of HPGe detectors. In low-background laboratories, HPGe detectors are optimised against background in the best possible configuration that allows for maximum sensitivity. Usually, the detectors are surrounded by thick layers of composite passive shields and, possibly, a muon veto system and/or a Rn abatement system may complete the setup heusser:1995 . In this regard, it is important to highlight that, in NAA, the effective background rate of a HPGe detector can be increased by the irradiated sample itself (correlated background). This happens when the sample matrix contains interfering elements that are activated because of their higher neutron cross sections and/or concentrations, as compared to the searched contaminants. As a result, a non-negligible background may be produced in the energy regions of the $\gamma$-lines of interest. This correlated background can significantly impair the NAA sensitivity if the interfering isotopes have decay times similar to or longer than the half-lives of the searched activation products. In such cases, possible solutions are the selection of the decay patterns of interest by means of $\beta-\gamma$ gespark or $\gamma-\gamma$ coincident systems (depending on the nuclear decay of the activated isotope), the discrimination of Compton events by pulse shape analysis, or the use of radiochemical treatments when the sample matrix allows it.
The aim of the analysis of the acquired $\gamma$ spectra is to get the number of counts $C$ (see Eq. 5) recorded in the sought-after peaks. If no peak is observed at the predicted energy, one can state that the maximum number of signal counts that could be emulated by a background fluctuation of a chosen significance level, expressed in numbers of Gaussian standard deviations ($n_{\sigma}$), is:
$C^{\text{max}}=n_{\sigma}\sqrt{B\,\Delta E\,T}$ (12)
@@
where $\Delta E$ is the energy range in which the peak is expected, usually set at 1.2 FWHM (Full Width at Half Maximum) standard-iso-2010 , $B$ is the background counting rate per unit energy, and $T$ the measurement time. The corresponding number of decays – to be used for absolute activity calculation – is given by:
$n_{\text{dec}}^{\text{max}}=\dfrac{C^{\text{max}}}{\varepsilon_{\gamma}\,I_{\gamma}\,f(\Delta E)}$ (13)
@@
where $f(\Delta E)$ is the fraction of signal events that fall in an energy window $\Delta E$ around the $\gamma$-ray energy. Then, by exploiting Eq. 4, one can set an upper limit on the activation rate $R$, which can be eventually converted into an upper limit for the contaminant concentration within the sample, using one of the methods outlined in the previous section.
@@ 3.1 Sample preparation
The presence of natural contaminants is ubiquitous in our environment, so the risk of contaminating a sample when aiming for ppt-level [^2] A contaminant concentration within a matrix is usually expressed in mBq/kg or g/g, i.e. grams of contaminant per gram of material, and its sub-multiples ppm $=10^{-6}$ g/g, ppb $=10^{-9}$ g/g, and ppt $=10^{-12}$ g/g). For convenience, we report the conversions between the two units: 1 mBq of 232Th activity per kg of material is equal to $2.5\times 10^{-10}$ g/g (or 0.25 ppb) of 232Th mass concentration in that material. Similarly, 1 mBq/kg of 238U corresponds to $8.1\times 10^{-11}$ g/g (or 81 ppt) of 238U, and 1 mBq/kg of 40K corresponds to $3.8\times 10^{-12}$ g/g (or 3.8 ppt) of 40K (or, equivalently, $3.1\times 10^{-8}$ g/g of natural K). sensitivities to their concentration within the sample itself is by no means negligible. Hence, the preparation of a sample and its insertion into the irradiation container must be followed with extreme care.
@@
In addition to liquid standards, we usually prepare also solid ones. These consist of certified strands of Al-Co (Co 0.1% wt.), a few mm in length and 1 mm in diameter, which are placed in each reactor channel used both for samples and liquid standards during NAA irradiation. This approach allows for cross-verification of results in case of any issues.
@@ 3.2 Irradiation at the TRIGA reactor in Pavia
As neutron source, we use the TRIGA Mark II nuclear research reactor of the University of Pavia, not far from the University of Milano-Bicocca. It is a pool-type reactor cooled and partly moderated by light water, with a maximum nominal power of 250 kW. Its cylindrical core ($\sim 46$ cm of diameter and $\sim 36$ cm in height) is surrounded by a 30-cm-thick radial graphite reflector and a 10-cm-thick axial graphite reflector at both the top and the bottom of each of the fuel elements. It is equipped with several irradiation facilities: among them, the Central Thimble, located at the center of the reactor core, and the Lazy Susan facility – a specimen rack in a circular well within the radial reflector – are the most suitable for our NAA analyses. The Lazy Susan facility is particularly attractive since it includes 40 irradiation channels and thus offers high flexibility in the number of samples to be irradiated. Each of these channels has a cylindrical shape with a diameter of $\sim 32$ mm; the PE irradiation vessels, therefore, also have a cylindrical shape, with an inner diameter of 23 mm and a useful height of $\sim 80$ mm. These are the dimensions that limit the size of the samples to be irradiated. The neutron fluxes within the Pavia TRIGA reactor are about $1.7\times 10^{13}$ n/cm2/s in the Central Thimble and about $2.2\times 10^{12}$ n/cm2/s in the Lazy Susan facility BayesianSpectrum . Typical operation times at full power for this TRIGA reactor are six hours per day. Consequently, our NAA campaigns are structured around six-hour irradiation periods for samples.
@@ 3.3 $\gamma$-spectroscopy measurements at Milano-Bicocca
After irradiation and once radiation protection measures allow, samples are promptly extracted from the reactor and sent to our Radioactivity Laboratory at the Department of Physics of the University of Milano-Bicocca for $\gamma$-spectroscopy. Timing is critical at this stage for achieving optimal sensitivity on 42K, given its half-life of only 12.36 hours. Typically, we can begin $\gamma$-counting in our laboratory about 4 hours after the reactor shutdown, assuming no unforeseen issues. However, interfering radioisotopes, such as 41Ar, 24Na, and 82Br in our case, may impede counting of relatively large samples after such short cooling times. In particular, the decay rate of 41Ar, which has a relatively short half-life of 109.6 minutes, may significantly increase the dead time of the HPGe detector. In such cases, we usually wait an additional 2 to 6 hours for 41Ar to decay.
Our Radioactivity Laboratory is equipped with five high-resolution HPGe detectors, with different relative efficiencies and background shieldings, along with one $\gamma-\gamma$ and one $\beta-\gamma$ low-background coincidence systems. Neutron irradiated samples for NAA analysis are usually counted on two of these detectors, named GeGEM and BEGe. GeGEM is a p-type coaxial HPGe detector by Ortec, with a relative efficiency of 30% and an energy resolution of 2.0 keV at the 1173 keV line of 60Co; it is surrounded by a composite shielding of 5 cm of copper and 12 cm of lead. BEGe is a planar-type broad energy HPGe by Canberra, with a relative efficiency of 50% and an energy resolution of 1.9 keV at the 1.33 MeV line of 60Co; it is surrounded by a composite shielding of 10 cm of copper and 15 cm of lead. The residual background rates of the two detectors in the main energy regions of the activation products of 41K, 232Th, and 238U are reported in Table 2. It is worth noting that these rates relate only to the constant background associated with the HPGe experimental setup and do not include the correlated time-dependent background originating from the sample after neutron irradiation.
@@ 3.4 Data analysis
To calculate the number of decays $n_{\text{dec}}$ in Eq. 5, it is crucial to determine the detector efficiency at full energy for each $\gamma$-peak. This is achieved through a Monte Carlo (MC) simulation, based on GEANT4 Geant4 , that accurately models the HPGe detector, its surrounding setup, and the geometry of the measurement, including the shape and position of the sample. To validate the simulation, the calculated efficiencies were iteratively compared with experimental values obtained from certified multi-$\gamma$ reference sources. This process continued until the agreement across the whole energy spectrum for all HPGe detectors in our laboratory was within 3%. We conservatively regard this deviation as the systematic uncertainty associated with efficiency evaluation in our NAA campaigns.

### v2

@@ Abstract
In the framework of physics experiments searching for rare events, the selection of extremely radiopure materials is a challenging task, as the signal of interest is often hidden by instrumental background. Neutron activation is a powerful technique to measure trace contaminants with high sensitivity but, to be properly applied, it requires a good characterization of neutron irradiation and $\gamma$-spectroscopy facilities. This paper presents the state-of-the-art workflow adopted by the Radioactivity Laboratory of the University of Milano-Bicocca for radiopurity screening of materials by neutron activation performed at the TRIGA reactor in Pavia. The ultimate sensitivity of the described workflow, in the absence of interfering activation products and without the application of radiochemical and/or active background rejection techniques, is $10^{-14}$ g/g for 40K ($10^{-10}$ g/g for elemental K), $10^{-12}$ g/g for 238U and of the order of $10^{-12}$ g/g for 232Th contaminations. Further details are here provided to address systematic uncertainties related to neutron irradiation that may bias results. To this aim, a dedicated neutron activation campaign was performed and the data were analyzed exploiting a Monte Carlo simulation model of the reactor and applying an unfolding technique to obtain a comprehensive characterization of the neutron flux in typical irradiation configurations. The results of this work provide a valuable benchmark for the application of neutron activation in future radiopurity screening campaigns.
@@ 1 Introduction
Experiments searching for rare particle physics events demand increasingly higher levels of radiopurity for the materials that make up the detectors heusser:1995 ; Laubenstein:2020rbe . The ultimate experimental sensitivity critically depends on the level of residual radioactive background in the energy region of interest, which could hamper the detection of the very faint signals being sought. Besides locating these experiments in deep underground laboratories to reduce cosmogenic-induced events Laubenstein:2004 , surrounding the detectors with massive shielding to attenuate environmental background Smith:2008 , and employing particle identification methods to discard spurious energy depositions Budjas:2009zu , it is crucial to construct the detector using materials with the lowest possible concentration of radioactive contaminants. In this respect, the assay techniques also require increasingly higher levels of sensitivity to evaluate the radiopurity of materials.
Typical contaminants include primordial radionuclides such as 40K, 232Th, 235U and 238U, as well as their decay chain daughters (for Th and U isotopes). Depending on the materials, a few cosmogenic and artificial long-lived radionuclides, such as 60Co and 137Cs, may also be of concern. The most widely used screening technique to study bulk contamination of materials is $\gamma$-spectroscopy Laubenstein:2017yjj , which has the significant advantage of being sensitive to most nuclides in natural decay chains and to many cosmogenic and anthropogenic radioisotopes. Alpha spectroscopy is typically used to evaluate the presence of 232Th, 238U, and their $\alpha$-decaying daughters (particularly radon progenies) on material surfaces Bunker:2020sxw . Complementary bulk contamination screening techniques include mass concentration measurement methods like Inductively Coupled Plasma Mass Spectrometry (ICP-MS) and Neutron Activation Analysis (NAA) Arpesella:2001iz ; Nisi:2009 . Although these techniques are limited to detecting only the parent nuclides, such as 232Th and 238U, without the ability to assess the concentrations of their daughters, they can achieve much higher analytical sensitivities compared to $\gamma$-spectroscopy.
@@ 2 Experimental methodology
The main step of NAA is neutron irradiation, usually at a research reactor. The contaminants to be measured are targets of neutron-induced reactions, which lead to the production of radionuclides. Subsequently, $\gamma$-spectroscopy measurements are performed to measure the induced activities and determine the concentrations of contaminants in the sample. High-Purity Germanium (HPGe) detectors are the most widely used $\gamma$-spectrometers, because of their excellent energy resolution.
@@
$^{238}\text{U}\xrightarrow{(n,\gamma)}\,^{239}\text{U}\xrightarrow[23.45\text{ m}]{\beta^{-}}\,^{239}\text{Np}\xrightarrow[2.356\text{ d}]{\beta^{-}}\,^{239}\text{Pu}$
$^{232}\text{Th}\xrightarrow{(n,\gamma)}\,^{233}\text{Th}\xrightarrow[22.3\text{ m}]{\beta^{-}}\,^{233}\text{Pa}\xrightarrow[26.97\text{ d}]{\beta^{-}}\,^{233}\text{U}$
@@
$n_{\text{dec}}=\int_{0}^{t_{\text{m}}}A_{0}\,e^{-\lambda t}dt=A_{0}\dfrac{1-e^{-\lambda t_{\text{m}}}}{\lambda}$ (3)
@@
where the superscript stdstd is used to label the parameters of the standard. It is worth noting here that if the standard has the same geometry and materials of the sample, the efficiency ratio $\varepsilon_{\gamma}/\varepsilon_{\gamma}^{std}$ can be assumed equal to 1, otherwise a Monte Carlo simulation is usually needed to determine this corrective factor with good accuracy.
@@
$\sigma_{\text{eff}}=\dfrac{\int\sigma(E)\,\varphi(E)\,dE}{\int\varphi(E)\,dE}$ (8)
and corresponds to the average cross section weighted on the flux spectrum. By calling $\Phi\equiv\int\varphi(E)\,dE$ the total flux intensity, Eq. 1 can be re-written as:
@@ 2.1 NAA sensitivity
After neutron irradiation, the samples are extracted from the reactor and prepared for $\gamma$-spectroscopy, typically by means of HPGe detectors. In low-background laboratories, HPGe detectors are optimised against background in the best possible configuration that allows for maximum sensitivity. Usually, the detectors are surrounded by thick layers of composite passive shields and, possibly, a muon veto system and/or a Rn abatement system may complete the setup heusser:1995 . In this regard, it is important to highlight that, in NAA, the effective background rate of a HPGe detector can be increased by the irradiated sample itself (correlated background). This happens when the sample matrix contains interfering elements that are activated because of their higher neutron cross sections and/or concentrations, as compared to the searched contaminants. As a result, a non-negligible background may be produced in the energy regions of the $\gamma$-lines of interest. This correlated background can significantly impair the NAA sensitivity if the interfering isotopes have decay times similar to or longer than the half-lives of the searched activation products. In such cases, possible solutions are the selection of the decay patterns of interest by means of $\beta-\gamma$ gespark ; BARRESI2025170322 or $\gamma-\gamma$ coincident systems (depending on the nuclear decay of the activated isotope), the discrimination of Compton events by pulse shape analysis BeGeCompton , or the use of radiochemical treatments when the sample matrix allows it.
The aim of the analysis of the acquired $\gamma$ spectra is to get the number of counts $C$ (see Eq. 5) recorded in the sought-after peaks. If no peak is observed at the predicted energy, one can calculate the detection limit for the signal counts $C^{*}$, applying the statistical method outlined by Currie in 1968 Currie1968 . This limit corresponds to the true signal that has only a 5% chance of fluctuating below the critical limit  [^2] The critical limit is defined as the value such that, if the result of a measurement exceeds it, one can conclude that a non-zero signal has been detected with a 95% confidence level.. Since in $\gamma$-spectroscopy it is usually possible to determine with good precision the background in the signal region, we calculate the Currie’s detection limit in the case of “well-known” background:
$C^{*}=2.71+3.29\,\sqrt{B\,t_{m}\,\Delta E}$ (12)
@@
where $B$ is the background counting rate per unit energy, $t_{m}$ the measurement time, and $\Delta E$ is the energy range in which the peak is expected, that we set at 1.2 FWHM (Full Width at Half Maximum) to maximize the signal-to-background ratio DEGEER2004151 ; standard-iso-2010 . The corresponding upper limit for the number of decays – to be used for absolute activity calculation – is given by:
$n_{\text{dec}}^{*}=\dfrac{C^{*}}{\varepsilon_{\gamma}\,I_{\gamma}\,f(\Delta E)}$ (13)
@@
where $f(\Delta E)$ is the fraction of signal events that fall in an energy window $\Delta E$ around the $\gamma$-ray energy ($f(\Delta E)=0.84$ when $\Delta E=1.2$ FWHM). Then, by exploiting Eq. 4, one can set an upper limit on the activation rate $R$, which can be eventually converted into an upper limit for the contaminant concentration within the sample, using one of the methods outlined in the previous section.
@@ 3.1 Sample preparation
The presence of natural contaminants is ubiquitous in our environment, so the risk of contaminating a sample when aiming for ppt-level [^3] A contaminant concentration within a matrix is usually expressed in mBq/kg or g/g, i.e. grams of contaminant per gram of material, and its sub-multiples ppm $10^{-6}$ g/g, ppb $10^{-9}$ g/g, and ppt $10^{-12}$ g/g). For convenience, we report the conversions between the two units: 1 mBq of 232Th activity per kg of material is equal to $2.5\times 10^{-10}$ g/g (or 0.25 ppb) of 232Th mass concentration in that material. Similarly, 1 mBq/kg of 238U corresponds to $8.1\times 10^{-11}$ g/g (or 81 ppt) of 238U, and 1 mBq/kg of 40K corresponds to $3.8\times 10^{-12}$ g/g (or 3.8 ppt) of 40K (or, equivalently, $3.1\times 10^{-8}$ g/g of natural K). sensitivities to their concentration within the sample itself is by no means negligible. Hence, the preparation of a sample and its insertion into the irradiation container must be followed with extreme care.
@@
In addition to liquid standards, we usually prepare also solid ones. These consist of certified strands of Al-Co (Co 0.5% wt.), a few mm in length and 1 mm in diameter, which are placed in each reactor channel used both for samples and liquid standards during NAA irradiation. This approach allows for cross-verification of results in case of any issues.
@@ 3.2 Irradiation at the TRIGA reactor in Pavia
As neutron source, we use the TRIGA Mark II nuclear research reactor of the University of Pavia, not far from the University of Milano-Bicocca. It is a pool-type reactor cooled and partly moderated by light water, with a maximum nominal power of 250 kW. Its cylindrical core ($\sim 46$ cm of diameter and $\sim 36$ cm in height) is surrounded by a 30-cm-thick radial graphite reflector and a 10-cm-thick axial graphite reflector at both the top and the bottom of each of the fuel elements. It is equipped with several irradiation facilities: among them, the Central Thimble, located at the center of the reactor core, and the Lazy Susan facility – a specimen rack in a circular well within the radial reflector – are the most suitable for our NAA analyses. More details about the geometry of the reactor and its irradiation facilities can be found in references AbsoluteFlux ; BayesianSpectrum ; FluxDistribution . The Lazy Susan facility is particularly attractive since it includes 40 irradiation channels and thus offers high flexibility in the number of samples to be irradiated. Each of these channels has a cylindrical shape with a diameter of $\sim 32$ mm; the PE irradiation vessels, therefore, also have a cylindrical shape, with an inner diameter of 23 mm and a useful height of $\sim 80$ mm. These are the dimensions that limit the size of the samples to be irradiated. The neutron fluxes within the Pavia TRIGA reactor are about $1.7\times 10^{13}$ n/cm2/s in the Central Thimble and about $2.2\times 10^{12}$ n/cm2/s in the Lazy Susan facility BayesianSpectrum . Typical operation times at full power for this TRIGA reactor are six hours per day. Consequently, our NAA campaigns are structured around six-hour irradiation periods for samples.
@@ 3.3 $\gamma$-spectroscopy measurements at Milano-Bicocca
After irradiation and once radiation protection measures allow, samples are promptly extracted from the reactor, put in a new container, and sent for $\gamma$-spectroscopy to our Radioactivity Laboratory at the third basement floor of the Department of Physics of the University of Milano-Bicocca. Timing is critical at this stage for achieving optimal sensitivity on 42K, given its half-life of only 12.36 hours. Typically, we can begin $\gamma$-counting in our laboratory about 4 hours after the reactor shutdown, assuming no unforeseen issues. However, interfering radioisotopes, such as 41Ar, 24Na, and 82Br in our case, may impede counting of relatively large samples after such short cooling times. In particular, the decay rate of 41Ar, which has a relatively short half-life of 109.6 minutes, may significantly increase the dead time of the HPGe detector. In such cases, we usually wait an additional 2 to 6 hours for 41Ar to decay.
Our Radioactivity Laboratory is equipped with five high-resolution HPGe detectors, with different relative efficiencies and background shieldings, along with one $\gamma-\gamma$ and one $\beta-\gamma$ low-background coincidence systems. Neutron irradiated samples for NAA analysis are usually counted on two of these detectors, named GeGEM and BEGe. GeGEM is a p-type coaxial HPGe detector by Ortec, with a relative efficiency of 30%, and a crystal mass of 815 g; it is surrounded by a composite shielding of 5 cm of copper and 12 cm of lead. BEGe is a planar-type broad energy HPGe by Canberra, with a relative efficiency of 50%, and a crystal mass of 802 g; it is surrounded by a composite shielding of 10 cm of copper and 15 cm of lead. Both detectors have an energy resolution of 1.9 keV at the 1.33 MeV line of 60Co. The residual background rates of the two detectors in the main energy regions of the activation products of 41K, 232Th, and 238U are reported in Table 2. It is worth noting that these rates relate only to the constant background associated with the HPGe experimental setup and do not include the correlated time-dependent background originating from the sample after neutron irradiation.
@@ 3.4 Data analysis
Figure 1: NAA detection limits achievable at the Radioactivity Laboratory of Milano-Bicocca (with the BeGE detector after 6 hours irradiation in the Lazy Susan facility of the TRIGA reactor of Pavia) for the concentration of the three natural contaminants in a generic plastic sample weighing 10 g, in the absence of interfering activation products. For a detailed description of the plots refer to the text. The uncertainties on neutron flux, cross sections and the other parameters entering in Eq. 14 are not represented here.
To calculate the number of decays $n_{\text{dec}}$ in Eq. 5, it is crucial to determine the detector efficiency at full energy for each $\gamma$-peak. This is achieved through a Monte Carlo (MC) simulation, based on GEANT4 Geant4 , that accurately models the HPGe detector, its surrounding setup, and the geometry of the measurement, including the shape and position of the sample. To validate the simulation, the calculated efficiencies were iteratively compared with experimental values obtained from certified multi-$\gamma$ reference sources. This process continued until the agreement across the whole energy spectrum for all HPGe detectors in our laboratory was within 5%. We conservatively regard this deviation as the systematic uncertainty associated with efficiency evaluation in our NAA campaigns.